If g is the inverse function f and f'(x) =1÷1+x^5 then g'(x) =

Grade:12

1 Answers

Khimraj
5 years ago
 
g is inverse of function f, then
f(g(x)) = x
differentiating wrt x
f’(g(x))*g’(x) = 1
So g’(x) = 1/f’(g(x)) = 1 + {g(x)}5
Hope it clears. If you like answer then please approve it.